OpenStudy (anonymous):

56 divided by 7/6 I got 336/7 is this right?

OpenStudy (tkhunny):

\(\dfrac{56}{\left(\dfrac{7}{6}\right)} = 56\cdot\dfrac{6}{7}\) This is where things went horribly astray. For some odd reason, you decided to do this: \(56\cdot\dfrac{6}{7} = \dfrac{56\cdot 6}{7} = \dfrac{336}{7}\) Then, you seem to have gotten stuck. Keep your eyes open for ways to proceed that do NOT make your life worse. \(56\cdot\dfrac{6}{7} = \dfrac{56}{7}\cdot 6 = 8\cdot 6 = 48\) I am NOT saying you managed an incorrect response. I am saying that you are going about it the hardest way possible. That may be fun for a while, but you will get bored and make more errors.

OpenStudy (anonymous):

ohh okay thank you so much